# MCITlib: Multimodal Continual Instruction Tuning Library and Benchmark

### v1

@@ 2 MCITlib Setup
LoRA-FT ICLR-22 58.03 77.63 44.39 67.40 61.77 58.22 76.89 61.24 68.57 -18.78
O-LoRA EMNLP-23 77.50 78.07 44.50 63.13 64.73 58.16 76.01 64.35 69.21 -13.99
MoELoRA NeurIPS-24 70.07 77.70 44.69 50.03 54.03 57.34 71.17 58.98 64.74 -14.63
ModalPrompt Arxiv-24 51.07 87.27 48.11 39.23 46.57 42.93 52.65 52.53 68.63 -0.15
ModalPrompt* Arxiv-24 74.43 92.00 55.92 44.27 53.97 43.67 60.76 60.71 62.59 -0.05
CL-MoE CVPR-25 66.33 77.00 44.78 51.87 53.53 57.42 71.46 58.49 63.94 -15.56
HiDE ACL-25 84.03 90.73 44.43 58.93 41.37 54.25 69.96 62.29 65.70 -9.20
SEFE ICML-25 80.83 78.00 47.01 69.63 65.83 57.92 75.98 66.54 70.25 -11.33
DISCO ICCV-25 87.43 93.07 46.96 68.13 65.70 56.69 75.87 69.66 72.71 -7.45
@@ 3 Experiment
LoRA-FT ICLR-22 69.65 41.59 25.43 40.88 87.45 64.98 53.00 58.52 -14.97
O-LoRA EMNLP-23 74.64 44.42 30.02 41.47 87.15 65.16 55.54 59.53 -12.03
MoELoRA NeurIPS-24 77.54 41.85 27.62 40.13 86.75 64.94 54.78 58.53 -12.71
ModalPrompt Arxiv-24 53.63 45.68 40.77 41.81 87.82 53.87 53.94 53.87 0.09
ModalPrompt* Arxiv-24 78.67 51.37 47.80 43.29 87.78 61.84 61.78 61.81 -0.07
CL-MoE CVPR-25 71.34 46.84 26.33 41.17 88.74 66.06 54.88 59.30 -13.97
HiDE ACL-25 74.31 48.95 33.21 38.54 81.55 60.77 55.31 57.04 -6.82
SEFE ICML-25 77.26 50.37 37.21 40.87 86.82 65.01 58.51 60.96 -8.13
DISCO ICCV-25 76.49 44.48 44.84 46.61 89.22 64.78 60.33 62.41 -5.57
@@
We report the numerical results of all methods on the two benchmarks in Table 2 and Table 3, while Figure 2 and Figure 3 illustrate their performance curves across training stages. First, we observe that in continual learning tasks for MLLMs, direct sequential fine-tuning does not lead to catastrophic forgetting as severely as in traditional class-incremental learning. This is reflected in the MFN results, where LoRA-FT surpasses the zero-shot baseline by 29.56% and 18.15% on the two benchmarks, respectively. This indicates that the strong inherent generalization ability of MLLMs can partially mitigate forgetting, though a substantial gap remains compared to the upper bound of individual task performance.

### v2

@@ 2 MCITlib Setup
LoRA-FT ICLR-22 58.03 77.63 44.39 67.40 61.77 58.22 76.89 61.24 76.55 -18.78
O-LoRA EMNLP-23 77.50 78.07 44.50 63.13 64.73 58.16 76.01 64.35 78.02 -13.99
MoELoRA NeurIPS-24 70.07 77.70 44.69 50.03 54.03 57.34 71.17 58.98 75.08 -14.63
ModalPrompt EMNLP-25 51.07 87.27 48.11 39.23 46.57 42.93 52.65 52.53 57.67 -0.15
ModalPrompt* EMNLP-25 74.43 92.00 55.92 44.27 53.97 43.67 60.76 60.71 70.57 -0.05
CL-MoE CVPR-25 66.33 77.00 44.78 51.87 53.53 57.42 71.46 58.49 74.19 -15.56
HiDE ACL-25 84.03 90.73 44.43 58.93 41.37 54.25 69.96 62.29 77.32 -9.20
SEFE ICML-25 80.83 78.00 47.01 69.63 65.83 57.92 75.98 66.54 78.76 -11.33
DISCO ICCV-25 87.43 93.07 46.96 68.13 65.70 56.69 75.87 69.66 81.60 -7.45
@@ 3 Experiment
LoRA-FT ICLR-22 69.65 41.59 25.43 40.88 87.45 64.98 53.00 61.13 -14.97
O-LoRA EMNLP-23 74.64 44.42 30.02 41.47 87.15 65.16 55.54 62.12 -12.03
MoELoRA NeurIPS-24 77.54 41.85 27.62 40.13 86.75 64.94 54.78 61.76 -12.71
ModalPrompt EMNLP-25 53.63 45.68 40.77 41.81 87.82 53.87 53.94 49.67 0.09
ModalPrompt* EMNLP-25 78.67 51.37 47.80 43.29 87.78 61.84 61.78 64.05 -0.07
CL-MoE CVPR-25 71.34 46.84 26.33 41.17 88.74 66.06 54.88 61.79 -13.97
HiDE ACL-25 74.31 48.95 33.21 38.54 81.55 60.77 55.31 60.68 -6.82
SEFE ICML-25 77.26 50.37 37.21 40.87 86.82 65.01 58.51 63.63 -8.13
DISCO ICCV-25 76.49 44.48 44.84 46.61 89.22 64.78 60.33 63.93 -5.57
@@
We present the numerical results of all methods on the two benchmarks in Table 2 and Table 3. Meanwhile, Figure 2 and Figure 3 illustrate the performance curves across training stages. The complete results matrix for each method is provided in Appendix A. First, we observe that in continual learning tasks for MLLMs, direct sequential fine-tuning does not lead to catastrophic forgetting as severely as in traditional class-incremental learning. This is reflected in the MFN results, where LoRA-FT surpasses the zero-shot baseline by 29.56% and 18.15% on the two benchmarks, respectively. This indicates that the strong inherent generalization ability of MLLMs can partially mitigate forgetting, though a substantial gap remains compared to the upper bound of individual task performance.
@@ 4 Conclusion
Supplementary Material
Appendix A Detailed Continual Learning Results
In this section, we show the detailed results for all methods at each stage of continual learning on LLaVA-1.5-7b.
ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
ImgNet-R 92.00
ArxivQA 90.63 91.03
VizWiz 74.00 91.4 61.9
IconQA 72.73 78.53 53.03 79.87
CLEVR 68.97 77.6 49.43 67.67 78.33
Flickr30k 58.03 77.63 44.39 67.4 61.77 58.22
Table 4: Result matrix of LoRA-FT on UCIT benchmark.
ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
ImgNet-R 91.40
ArxivQA 90.07 94.73
VizWiz 81.50 91.70 61.59
IconQA 82.4 77.73 54.06 76.63
CLEVR 81.07 77.87 50.62 62.00 73.53
Flickr30k 77.50 78.07 44.50 63.13 64.73 58.16
Table 5: Result matrix of O-LoRA on UCIT benchmark.
ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
ImgNet-R 91.23
ArxivQA 90.07 92.80
VizWiz 80.30 92.03 60.22
IconQA 80.53 77.20 51.19 53.73
CLEVR 79.37 77.40 48.68 51.30 71.67
Flickr30k 70.07 77.70 44.69 50.03 54.03 57.34
Table 6: Result matrix of MoELoRA on UCIT benchmark.
ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
ImgNet-R 51.10
ArxivQA 51.03 87.37
VizWiz 51.10 87.37 48.14
IconQA 51.20 87.27 48.03 39.30
CLEVR 51.17 87.23 47.96 39.40 47.07
Flickr30k 51.07 87.27 48.11 39.23 46.57 42.93
Table 7: Result matrix of ModalPrompt on UCIT benchmark.
ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
ImgNet-R 74.57
ArxivQA 74.73 91.83
VizWiz 74.53 91.93 55.76
IconQA 74.53 91.90 55.97 44.20
CLEVR 74.50 91.83 55.75 44.23 54.50
Flickr30k 74.43 92.00 55.92 44.27 53.97 43.67
Table 8: Result matrix of ModalPrompt* on UCIT benchmark.
ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
ImgNet-R 91.30
ArxivQA 87.40 92.27
VizWiz 76.23 90.53 60.18
IconQA 77.33 76.20 51.09 54.33
CLEVR 75.73 75.70 48.27 52.80 73.23
Flickr30k 66.33 77.00 44.78 51.87 53.53 57.42
Table 9: Result matrix of CL-MoE on UCIT benchmark.
ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
ImgNet-R 90.87
ArxivQA 88.83 92.27
VizWiz 88.17 91.90 54.46
IconQA 87.00 89.70 46.51 73.13
CLEVR 86.23 90.07 44.28 64.37 54.77
Flickr30k 84.03 90.73 44.43 58.93 41.37 54.25
Table 10: Result matrix of HiDe on UCIT benchmark.
ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
ImgNet-R 90.80
ArxivQA 90.23 91.23
VizWiz 83.60 89.73 61.24
IconQA 84.17 77.20 56.36 78.93
CLEVR 82.90 77.13 53.54 71.43 75.73
Flickr30k 80.83 78.00 47.01 69.63 65.83 57.92
Table 11: Result matrix of SEFE on UCIT benchmark.
ImgNet-R ArxivQA VizWiz IconQA CLEVR Flickr30k
ImgNet-R 90.97
ArxivQA 90.07 91.77
VizWiz 89.40 92.53 61.39
IconQA 88.50 93.97 58.61 78.00
CLEVR 88.03 93.40 57.11 70.93 76.40
Flickr30k 87.43 93.07 46.96 68.13 65.70 56.69
Table 12: Result matrix of DISCO on UCIT benchmark.
RS Med AD Sci Fin
RS 78.32
Med 74.68 57.53
AD 68.93 47.19 52.15
Sci 75.12 45.56 38.46 49.44
Fin 69.65 41.59 25.43 40.88 87.45
Table 13: Result matrix of LoRA-FT on DCL benchmark.
RS Med AD Sci Fin
RS 79.25
Med 74.05 56.52
AD 76.06 43.71 52.32
Sci 76.60 44.87 40.57 50.58
Fin 74.64 44.42 30.02 41.47 87.15
Table 14: Result matrix of O-LoRA on DCL benchmark.
RS Med AD Sci Fin
RS 79.09
Med 74.78 58.73
AD 77.69 43.72 51.47
Sci 76.87 43.79 32.81 48.67
Fin 77.54 41.85 27.62 40.13 86.75
Table 15: Result matrix of MoELoRA on DCL benchmark.
RS Med AD Sci Fin
RS 53.12
Med 53.02 45.82
AD 53.14 45.77 40.78
Sci 53.03 45.65 40.78 41.83
Fin 53.63 45.68 40.77 41.81 87.82
Table 16: Result matrix of ModalPrompt on DCL benchmark.
RS Med AD Sci Fin
RS 78.76
Med 78.57 51.46
AD 78.61 51.59 47.92
Sci 78.59 51.50 47.90 43.28
Fin 78.67 51.37 47.80 43.29 87.78
Table 17: Result matrix of ModalPrompt* on DCL benchmark.
RS Med AD Sci Fin
RS 78.76
Med 78.57 51.46
AD 78.61 51.59 47.92
Sci 78.59 51.50 47.90 43.28
Fin 78.67 51.37 47.80 43.29 87.78
Table 18: Result matrix of ModalPrompt* on DCL benchmark.
RS Med AD Sci Fin
RS 79.08
Med 73.48 60.56
AD 72.61 44.42 51.62
Sci 71.02 48.04 37.70 50.28
Fin 71.34 46.84 26.33 41.17 88.74
Table 19: Result matrix of CL-MoE on DCL benchmark.
RS Med AD Sci Fin
RS 78.14
Med 74.26 58.05
AD 74.90 42.94 39.65
Sci 75.43 44.91 38.33 46.44
Fin 74.31 48.95 33.21 38.54 81.55
Table 20: Result matrix of HiDe on DCL benchmark.
RS Med AD Sci Fin
RS 78.27
Med 76.32 58.42
AD 77.22 49.13 52.49
Sci 77.83 47.7 43.01 49.04
Fin 77.26 50.37 37.21 40.87 86.82
Table 21: Result matrix of SEFE on DCL benchmark.
RS Med AD Sci Fin
RS 78.57
Med 75.80 52.36
AD 76.37 49.78 53.04
Sci 76.11 45.76 49.26 49.23
Fin 76.03 45.20 43.79 42.33 88.95
Table 22: Result matrix of DISCO on DCL benchmark.